# ChatGPT in Data Visualization Education: A Student Perspective

### v1

@@ Abstract
Unlike traditional educational chatbots that rely on pre-programmed responses, large-language model-driven chatbots, such as ChatGPT, demonstrate remarkable versatility and have the potential to serve as a dynamic resource for addressing student needs from understanding advanced concepts to solving complex problems. This work explores the impact of such technology on student learning in an interdisciplinary, project-oriented data visualization course. Throughout the semester, students engaged with ChatGPT across four distinct projects, including data visualizations and implementing them using a variety of tools including Tableau, D3, and Vega-lite. We collected conversation logs and reflection surveys from the students after each assignment. In addition, we conducted interviews with selected students to gain deeper insights into their overall experiences with ChatGPT. Our analysis examined the advantages and barriers of using ChatGPT, students’ querying behavior, the types of assistance sought, and its impact on assignment outcomes and engagement. Based on the findings, we discuss design considerations for an educational solution that goes beyond the basic interface of ChatGPT, specifically tailored for data visualization education.
@@
ChatGPT, data visualization, education, project-based learning
@@ I Introduction
  • What are the advantages and barriers of ChatGPT use in data visualization tasks? $\rightarrow$ We look into survey responses and interview results across different assignments.
@@
  • What is the impact of ChatGPT on their learning outcomes and engagement? $\rightarrow$ We investigate how various factors, including the volume of conversations and the polarity of attitudes, relate to course outcomes.
@@
Upon analyzing the conversation logs, we identified three high-level querying themes in order of frequency: solution-oriented queries (e.g., “help me create HTML”), abstract and open-ended queries (e.g., “make my website prettier”), and interpersonal queries (e.g., “margins not applying!”). We initially hypothesized that more frequent and positive use of ChatGPT, particularly solution-oriented queries, might lead to better grades. However, when analyzing the relationship between the query themes and student grades, we did not observe any significant impacts. Similarly, our examination of the correlation between the number of interactions, query lengths, and grades also revealed no significant effects.
Based on the lessons learned from our study, we discuss design considerations for a future educational system tailored to data visualization education. These considerations include promoting design-oriented queries, supporting diverse inputs and outputs for fluid interactions, and evolving beyond merely being a tool for problem-solving. Our contribution lies in providing an improved understanding of how students use ChatGPT in design-oriented, project-based courses.
@@ II Background & Related Work
The integration of artificial intelligence in educational settings has significantly evolved over the past decades, transitioning from early intelligent tutoring systems that mimicked human-tutor interactions based on rule-based reasoning, to more advanced systems employing machine learning and natural language processing [22, 38]. Example tutoring systems include Cognitive Tutors, which uses model-tracing to provide step-by-step coaching [6], AutoTutor, which engages students in conversational interactions [16], and ASSISTments, which blends tutoring assistance and assessment reporting [18].
Intelligent tutors, also known as AI chatbots, are used across various educational domains, serving roles such as subject tutors in science and mathematics, administrative assistants for course scheduling, and support for special education needs related to cultural diversity and accessibility [42, 21]. Several authoring tools have been developed to facilitate the creation of these intelligent tutors, including the Cognitive Tutor Authoring Tools [5], ASSISTments Builder [44], and AutoTutor tools [39]. More recently, general-purpose cloud-based chatbot platforms such as Dialogflow [1], Wit.ai [3], and Rasa [2] have simplified the process of building custom conversational agents [13]. These chatbots adopt various roles, including teachers and learning companions, and exhibit diverse interaction styles ranging from user-driven to machine-driven conversations [32].
@@
Recently, chatbots based on large language models (LLMs), notably ChatGPT, have garnered widespread attention across society. These chatbots, trained on extensive datasets from the web, display a high degree of versatility in performing diverse tasks across various disciplines, including education. In response to this significant advancement in AI, there has been a surge in scholarly inquiries and investigations aimed at exploring the potential of these technologies in educational settings. Researchers discuss the potential benefits and threats based on prior literature [27, 33, 31] and reactions on social media [47, 36]. Generally, they highlight similar advantages to previous generations, such as serving as personal learning aids or instructional assistant tools, while also illustrating more diverse scenarios due to the versatility of LLMs, such as general content generation and problem-solving [36, 33]. On the other hand, there is increased concern regarding potentially exacerbating drawbacks, such as biases and errors, as well as a growing reliance on AI [31, 27]. Researchers argue that educators need to prepare for changes in educational practices, including developing new pedagogical approaches and addressing issues of academic integrity [8, 47].
@@
Despite the relatively short history of LLM-based chatbots, numerous studies have explored the perspectives of teachers and students. Yu discusses potential coping strategies for teachers, including developing collaborative pedagogy with AI, nurturing responsible AI engagement, and cultivating emotional intelligence [52]. Herft offers guidance on various use cases for prompting, ranging from creating lesson plans and teaching strategies to generating formative assessment questions [19]. Ghimire et al. conducted surveys and interviews with educators to gauge their viewpoints, revealing an increasing awareness and positive perceptions, particularly among computer science educators [15]. Jeon and Lee identified complementary roles for ChatGPT and human teachers [23]. Other researchers have examined ChatGPT’s capabilities in recommending teaching instructions [24], generating new assignments [25], and completing existing assignments [12, 46, 41, 26].
@@
Numerous studies have found that students generally hold positive attitudes and experiences with ChatGPT [45, 11, 10]. Students use it for a variety of tasks including writing essays, understanding complex concepts, exploring alternative solutions, and enhancing programming skills [51, 10]. Another set of studies explored how students use LLM-based code generators and assistants in practical courses, identifying various coding approaches and usability considerations [29, 35, 43, 28]. In controlled experiments, Vaithilingam found that using Copilot led to more frequent failures compared to using Intellisense [48], while Mousa and Veilleux discovered that students who did not use ChatGPT in Mathematics performed better than those who did [37]. Hou et al.’s study indicates that Gen AI has not yet surpassed traditional help resources, and that help-seeking with Gen AI is a skill that requires development [20].
@@ III Methodology
In this section, we explain the tasks & procedures and participant demographics used to explore the three research goals described in the introduction.
@@ III-A Procedures
Participants were given four assignments during the course. For each assignment, students were instructed to log their conversations with ChatGPT using a dedicated chat instead of creating new chats every time they used it. They were asked to submit the log in the markdown format using the Save ChatGPT Chrome extension. Following completion of each assignment, they were required to fill out a reflection survey regarding their experience.
Upon conclusion of the course, students filled out the final post-course survey. Additionally, selected students participated in follow-up interviews, which required separate consent forms and were conducted using Zoom, with audio recordings saved for later analysis. During the interviews, participants were queried about their experience using ChatGPT for the assignments, with questions tailored to their survey responses. The interviews took place between December 15, 2023, and January 4, 2024. The interviews varied in length, ranging from 17 to 38 minutes, with a median length of 26 minutes.
Participation in the study was voluntary. Participants in the study received an additional 0.5% extra credit for each assignment, with the total opportunity for extra credit being 2%. Students who chose not to participate in the study were offered an alternative method to earn the same extra credit by completing a separate assignment where they were tasked with creating an original visualization. Those who completed the interviews received a \$25 Amazon gift card as compensation.
@@ III-B1 Survey participants
Regarding data visualization experience, more than half of the participants (15/26) had not been introduced to or worked on data visualization prior to this course. Most identified themselves as beginners in data visualization (19/26), with 6 at an intermediate level, and only 1 participant self-identified as advanced. The tools used for visualization varied, with Excel or Google Sheets being the most common (23/26), followed by R or Python for statistical visualization (16/26), and a few used more specialized software like Tableau (2/26), PowerBI (3/26), and Vega-lite (2/26). The majority had engaged in creating visualizations for academic assignments or projects (20/26), while 15 had done so for work-related purposes, three had no prior experience in creating visualizations before this course, and three had created visualizations for personal projects or for fun.
@@ III-B2 Interview Participants
Table I: Demographics of interview participants: grades are in percentages, ChatSize represents the number of interactions with ChatGPT, exam scores are from an in-class quiz on lecture material, and final grades include extra credit.
@@ III-C3 Interactive Dashboard using D3
In the third assignment, students are tasked with constructing multiple coordinated interactive visualizations using D3. They were provided with the same datasets as the second assignment. Technical requirements were in place including at least three visualizations with at least two different chart types and coordinating interactions. The goal is to practice web-based visualization implementation skills.
@@ IV-A Survey & Interview Analysis
Our final dataset is derived from 26 participants and includes their survey responses to the four assignments, the end-of-course survey, and conversational logs. Additionally, we have interview transcripts from 14 of the original 26 participants.
For analyzing surveys, we computed descriptive statistics. For interviews, two researchers performed an open coding of the transcribed interviews. Initial themes were derived from initial questions such as asking about strengths and weaknesses as well as feedback & improvements. The researchers iteratively reviewed the transcribed responses in multiple meetings and collaboratively recategorized them into newly emerging themes, primarily based on advantages ( +) and barriers ( -).
@@ IV-B Conversational Log Analysis
We gathered 3,773 user queries we collected from 26 participants. Since the original query data is often too long, we manually processed user queries by replacing parts that are not essential for understanding the overall themes of the query dataset. Specifically, we replaced the Javascript, Python, HTML code into $<$CODE$>$, csv or tsv format data into $<$DATA$>$, https link to $<$LINK$>$, evaluation rubrics and excessive textual explanation into $<$INFO$>$, and error message to $<$ERROR_MSG$>$.
To perform thematic analysis more efficiently, we leveraged the automatic qualitative coding approach, suggested by Hämäläinen et al. [17]. The method simulates the overall process of humans doing the thematic analysis by leveraging large language models. We generated representative codes for each query by feeding the following prompt to GPT-4-turbo by OpenAI:
@@
We limited the number of codes to a maximum of five to prevent overfitting. After manually filtering out irrelevant and erroneous codes, we obtained 16,017 valid codes from an initial 18,865. We then preserved 4,043 unique codes and vectorized them using OpenAI embeddings  [^1] https://platform.openai.com/docs/models/embeddings, reducing their dimensions from 1,536 to five to better capture the semantic relationships between the vectors. Subsequently, we used the HDBSCAN algorithm to cluster these vectors into a small number of semantically relevant classes. We further consolidated these clusters manually into fewer, higher-level clusters to extract broader themes. We then examined the distribution of the themes among different users and assignments.
@@ V-A Final Course Survey Result
ChatGPT was found most valuable for the “Interactive dashboard using D3” project (54%), while it was deemed least useful for “Exploration and exposition with Tableau” (62%). Key strengths of ChatGPT identified by participants included its assistance in “Code Syntax and Debugging” (89%), “Being Available 24/7” (85%), and its ability to “Retrieve Information Quickly” (65%). On the other hand, notable challenges faced by users were the “Inaccuracy of Information” (78%), the tendency to develop an ”Over-reliance on AI for Problem-Solving” (59%), and the “Difficulty of Phrasing Questions” for ChatGPT (44%).
@@ V-C Post-Course Interview Results
V-C2 - Not Used Enough for Visualization Designs
Although rubrics were in place for evaluating the quality of visualization designs, students did not utilize ChatGPT as much for coding assistance (P2, P6, P10, P11, P12, P14). For instance, P10 said “I think for a lot of my projects, the main reason I use ChatGPT was to understand the code better and to have to be able to implement the code from class into our projects”, while P14 shared “But anything we kind of learned in class and anything about design, or anything regarding creativity and stuff like that, I just revisited lecture slides for that.”
On the other hand, several participants mentioned use cases for design knowledge. They mentioned chart suggestions as a useful feature (P3, P5 P7), such as showing a snapshot of the data and asking what chart would fit the data (P5, P7) or asking for an alternative representation to avoid the overuse of a bar chart (P9). Others also used ChatGPT to justify their design choices and identify the pros and cons of a chart (P3)—which was part of the first assignment requirement—asking for feedback for color schemes (P5, P7), and craft text narratives given charts (P3, P13). P3 gave a reason for the use case of design rationale generation, saying “I could think of some of my own, but I thought ChatGPT articulated them better because it knew more like design choice words than I did at the time.”
@@
Most participants experienced the occurrence of mistakes in the code generated by ChatGPT or response outputs that did not meet their expectations (P1, P2, P3, P6, P11, P13, P14). For instance, P2 expressed frustration, stating, “it saved me a good amount of time writing all the code. But sometimes it would make a little mistake, which would be annoying, because now I’d have to like, a lot of effort to fix it.” Often participants observed ChatGPT pretend to understand user questions but keep generating incorrect responses, increasing the level of frustration (P1, P3, P13), as P13 described, “[…] it cannot solve my problems, but it still pretends that I can, I will still believe it. And that just makes my life harder.” Oftentimes, participants ended up asking repetitive follow-up questions (P6, P11) as P11 shared “going into an endless loop with ChatGPT telling it that it’s wrong. And then having to like, in the end, figure stuff out myself anyway.”
@@
Several participants expressed worries about becoming overly dependent on ChatGPT (P1, P3, P4, P8, P11, P12, P13). P1 cautioned about the “over-reliance on AI,” highlighting a tendency to “let it do it for you”. P6 and P7 also mirrored these concerns, indicating a habitual turn to ChatGPT as the first response to coding problems. P14 felt that ChatGPT was “most helpful if I already understood the concept, but just didn’t really want to spend the time implementing it myself.”, implying a trade-off between efficiency and educational depth.
@@
V-D5 - Not Necessarily for Enhancing Engagement or Curiosity

### v2

@@ Abstract
Unlike traditional educational chatbots that rely on pre-programmed responses, large-language model-driven chatbots, such as ChatGPT, demonstrate remarkable versatility to serve as a dynamic resource for addressing student needs from understanding advanced concepts to solving complex problems. This work explores the impact of such technology on student learning in an interdisciplinary, project-oriented data visualization course. Throughout the semester, students engaged with ChatGPT across four distinct projects, designing and implementing data visualizations using a variety of tools such as Tableau, D3, and Vega-lite. We collected conversation logs and reflection surveys after each assignment and conducted interviews with selected students to gain deeper insights into their experiences with ChatGPT. Our analysis examined the advantages and barriers of using ChatGPT, students’ querying behavior, the types of assistance sought, and its impact on assignment outcomes and engagement. We discuss design considerations for an educational solution tailored for data visualization education, extending beyond ChatGPT’s basic interface.
@@
ChatGPT, large language model, data visualization, education, project-based learning
@@ I Introduction
  • What are the challenges of ChatGPT use in data visualization tasks? $\rightarrow$ We look into survey responses and interview results across different assignments.
@@
  • What is the impact of ChatGPT on their learning outcomes and engagement? $\rightarrow$ We examine how factors like conversation volume and attitude relate to course grades.
@@
Upon analyzing the conversation logs, we identified three high-level querying themes in order of frequency: solution-oriented queries (seeking a specific outcome with a clear goal, such as “help me create HTML”), abstract and open-ended queries (exploring options without a definitive outcome, such as “make my website prettier”), and interpersonal queries (expressing emotional feedback such as, “margins not applying!”). We initially hypothesized that more determined and frequent use of ChatGPT might lead to better grades. However, when analyzing the relationship between the query themes and student grades, we did not observe any significant impacts. Similarly, our examination of the correlation among the number of interactions, query lengths, and grades also revealed no significant effects.
Based on the lessons learned from our study, we discuss design considerations for a future educational system tailored to data visualization education. These considerations include promoting design-oriented queries, supporting diverse inputs and outputs for fluid interactions, and evolving beyond merely being a tool for chart implementation. Our contribution lies in providing an improved understanding of how students use ChatGPT in a design-oriented, project-based course.
@@ II Background & Related Work
The integration of artificial intelligence in educational settings has evolved over the past decades, transitioning from rule-based tutoring systems that mimicked human-tutor interactions to more advanced systems employing machine learning and natural language processing [24, 41]. Examples include Cognitive Tutors, which uses model-tracing to provide step-by-step coaching [6], AutoTutor, which engages students in conversational interactions [18], and ASSISTments, which blends tutoring assistance and assessment reporting [20].
Intelligent tutors are used across various educational domains, serving roles such as subject tutors in science and mathematics, administrative assistants for course scheduling, and support for special education needs related to cultural diversity and accessibility [45, 23]. Several authoring tools have been developed to facilitate the creation of these intelligent tutors, including the Cognitive Tutor Authoring Tools [5], ASSISTments Builder [47], and AutoTutor tools [42]. More recently, general-purpose cloud-based chatbot platforms such as Dialogflow [1], Wit.ai [3], and Rasa [2] have simplified the process of building custom conversational agents [15]. These chatbots adopt various roles, including teachers and learning companions, and exhibit diverse interaction styles ranging from user-driven to machine-driven conversations [34].
@@
Recently, chatbots based on large language models, notably ChatGPT, have garnered widespread attention across society. These chatbots, trained on extensive datasets from the web, display a high degree of versatility in performing diverse tasks across various disciplines, including education. This AI advancement has sparked numerous scholarly inquiries and investigations into its educational potential. Researchers discuss the potential benefits and threats based on prior literature [29, 35, 33] and reactions on social media [52, 39]. Generally, they highlight similar advantages to previous generations, such as serving as personal learning aids or instructional assistant tools, while also illustrating more diverse scenarios due to the versatility of LLMs, such as general content generation and problem-solving [39, 35]. On the other hand, there is increased concern regarding potentially exacerbating drawbacks, such as biases and errors, as well as a growing reliance on AI [33, 29]. Researchers argue that educators need to prepare for changes in educational practices, including developing new pedagogical approaches and addressing issues of academic integrity [9, 52].
@@
Despite the relatively short history of LLM-based chatbots, several studies have explored the perspectives of teachers and students. Yu discusses potential coping strategies for teachers, including developing collaborative pedagogy with AI, nurturing responsible AI engagement, and cultivating emotional intelligence [58]. Herft offers guidance on various use cases for prompting, ranging from creating lesson plans and teaching strategies to generating formative assessment questions [21]. Others conducted interviews with educators to gauge their viewpoints, revealing an increasing awareness, positive and negative perceptions, and ethical concerns [17, 55]. Jeon and Lee identified complementary roles for ChatGPT and human teachers [25]. Other researchers have examined ChatGPT’s capabilities in recommending teaching instructions [26], generating new assignments [27], and completing existing assignments [14, 50, 44, 28, 55].
@@
Other studies have found that students generally hold positive attitudes and experiences with ChatGPT [48, 13, 11]. Students use it for a variety of tasks including writing essays, understanding complex concepts, exploring alternative solutions, and enhancing programming skills [57, 11]. Another set of studies explored how students use LLM-based code generators and assistants in practical courses, identifying various coding approaches and usability considerations [31, 38, 46, 30]. In controlled experiments, Vaithilingam found that using Copilot led to more frequent failures compared to using Intellisense [53], while Mousa and Veilleux discovered that students who did not use ChatGPT in Mathematics performed better than those who did [40]. Hou et al.’s study indicates that Gen AI has not yet surpassed traditional help resources, and that help-seeking with Gen AI is a skill that requires development [22].
@@ III Methodology
This section describes the procedures, participants, and tasks used to address the research goals outlined earlier.
@@ III-A Procedures
Participants completed four assignments during the course (see Section III-C Tasks). For each assignment, students were instructed to log their conversations with ChatGPT using a dedicated chat and asked to submit the log in the markdown format using the Save ChatGPT Chrome extension. After each assignment, they were required to fill out a reflection survey regarding their experience.
Upon course completion, students completed a final post-course survey. Selected students also participated in follow-up Zoom interviews, which required separate consent forms. Audio recordings were saved for later analysis. Interviewees were asked about their ChatGPT experience, with questions based on their survey responses. The interviews took place between December 15, 2023, and January 4, 2024. The interviews varied in length, ranging from 17 to 38 minutes, with a median length of 26 minutes.
Participation in the study was voluntary. Students received 0.5% extra credit per assignment, up to 2% total. Students who chose not to participate could earn equivalent credit by completing a separate assignment where they were tasked with creating an original visualization. Interviewees received a \$25 Amazon gift card as compensation.
@@ III-B1 Survey participants
Regarding data visualization experience, more than half of the participants (15/26) had not been introduced to or worked on data visualization prior to this course. Most identified themselves as beginners in data visualization (19/26), with 6 at an intermediate level, and only 1 participant self-identified as advanced. The tools used for visualization varied, with Excel or Google Sheets being the most common (23/26), followed by R or Python for statistical visualization (16/26), and a few used more specialized software like PowerBI (3/26), Tableau (2/26), and Vega-lite (2/26). The majority had engaged in creating visualizations for academic assignments or projects (20/26), while 15 had done so for work-related purposes, three had no prior experience in creating visualizations before this course, and three had created visualizations for personal projects or for fun.
@@ III-B2 Interview Participants
Table I: Demographics of interview participants: grades are in percentages, ChatSize represents the number of turn-taking interactions with ChatGPT, exam scores are from an in-class quiz on lecture material, and final grades include extra credit.
@@ III-C3 Interactive Dashboard using D3
In the third assignment, students are tasked with constructing multiple coordinated interactive visualizations using D3. They were provided with the same datasets as the second assignment. Technical requirements were in place including at least three visualizations with at least two different chart types and coordinating interactions. The goal is to practice web-based interactive visualization implementation skills.
@@ IV Data & Analysis Methods
Our final dataset comprises survey responses, conversational logs, and end-of-course surveys from 26 participants, plus interview transcripts from 14 of these participants.
@@ IV-A Survey & Interview Analysis
For analyzing surveys, we primarily used frequency analysis to compute percentage distributions of survey responses. For interviews, two researchers conducted open coding of the transcripts. They organized quotes in a spreadsheet based on themes and categories, initially derived from interview questions (e.g., strengths, weaknesses, feedback, and improvements). Through multiple collaborative meetings, they refined and recategorized responses into emerging themes, resolving conflicts as they arose. The final themes, categorized as advantages ( +) or barriers ( -), include coding assistance, accessibility, serendipitous learning, and task-dependent effectiveness. These themes are directly presented in Section V-C, with participant IDs for transparency.
@@ IV-B Conversational Log Analysis
We gathered 3,773 user queries from 26 participants. Since the original query data is often too long, we manually processed user queries by replacing parts that are not essential for understanding the overall themes of the query dataset. Specifically, we replaced the Javascript, Python, HTML code into $<$CODE$>$, csv or tsv format data into $<$DATA$>$, https link to $<$LINK$>$, evaluation rubrics and excessive textual explanation into $<$INFO$>$, and error message to $<$ERROR_MSG$>$.
To perform thematic analysis more efficiently, we employed an automatic qualitative coding approach leveraging large-language models, following the recent approach  [19]. The method simulates the overall process of humans doing the thematic analysis. First, we generated representative codes for each query by feeding the following prompt to GPT-4-turbo by OpenAI:
@@
We limited the number of codes to a maximum of five per query to prevent overfitting. After manually examining the codes for quality assurance and filtering out irrelevant and erroneous entries, we reduced the initial 18,865 codes to 16,017 valid codes. We then consolidated overlapping codes, resulting in 4,043 unique codes. To capture broader themes from these codes, we applied a clustering algorithm. First, we computed semantic embedding vectors of the codes using OpenAI embeddings  [^1] https://platform.openai.com/docs/models/embeddings, and then reduced the dimensionality of these vectors using UMAP. Next, we applied HDBSCAN to combine similar codes into semantically relevant topics. We evaluated topic prevalence by counting the frequencies of the code clusters (i.e., the number of user queries assigned at least one code from each cluster). Through manual examination, we further consolidated these clusters into three higher-level themes, as described in Section V-E. We also examined the distribution of these themes among different users and assignments to gain insights into their relative importance and occurrence patterns.
@@ V-A Final Course Survey Result
ChatGPT was found most valuable for the “Interactive dashboard using D3” project (54%), while it was deemed least useful for “Exploration and exposition with Tableau” (62%). Key strengths of ChatGPT identified by participants included its assistance in “Code Syntax and Debugging” (89%), “Being Available 24/7” (85%), and its ability to “Retrieve Information Quickly” (65%). On the other hand, notable challenges faced by users were the “Inaccuracy of Information” (78%), the tendency to develop an ”Over-reliance on AI for Problem-Solving” (59%), and the “Difficulty of Phrasing Questions” for ChatGPT (44%). These findings corroborate prior survey results that report students’ use of ChatGPT, both in general and in programming courses. [48, 13, 11, 57].
@@ V-C Post-Course Interview Results
We categorize findings as advantages or barriers, but their impact varies among participants. Some students may overcome barriers that others find challenging. We report the main findings along with these exceptional cases as examples.
@@
V-C2 - Limited Use for Visualization Design Guidance
Although rubrics were in place for evaluating the quality of visualization designs, students did not utilize ChatGPT as much for design assistance (P2, P6, P10-P12, P14). For instance, P10 said “I think for a lot of my projects, the main reason I use ChatGPT was to understand the code better and to have to be able to implement the code from class into our projects”, while P14 shared “But anything we kind of learned in class and anything about design, or anything regarding creativity and stuff like that, I just revisited lecture slides for that.”
Although limited compared to what they learned in classes (e.g., choosing effective encoding choices based on perception, tasks, and data types), participants mentioned several basic use cases for seeking design knowledge. For instance, they mentioned chart suggestions as a useful feature (P3, P5 P7), such as showing a snapshot of the data and asking what chart would fit the data (P5, P7) or asking for an alternative representation to avoid the overuse of a bar chart (P9). Others also used ChatGPT to justify their design choices and identify the pros and cons of a chart (P3)—which was part of the first assignment requirement—asking for feedback for color schemes (P5, P7), and craft text narratives given charts (P3, P13). P3 gave a reason for the use case of design rationale generation, saying “I could think of some of my own, but I thought ChatGPT articulated them better because it knew more like design choice words than I did at the time.”
@@
Most participants experienced the occurrence of mistakes in the code generated by ChatGPT or response outputs that did not meet their expectations (P1-P3, P6, P11, P13, P14). For instance, P2 expressed frustration, stating, “it saved me a good amount of time writing all the code. But sometimes it would make a little mistake, which would be annoying, because now I’d have to like, a lot of effort to fix it.” Often participants observed ChatGPT pretend to understand user questions but keep generating incorrect responses, increasing the level of frustration (P1, P3, P13), as P13 described, “[…] it cannot solve my problems, but it still pretends that I can, I will still believe it. And that just makes my life harder.” Oftentimes, participants ended up asking repetitive follow-up questions (P6, P11) as P11 shared “going into an endless loop with ChatGPT telling it that it’s wrong. And then having to like, in the end, figure stuff out myself anyway.”
@@
Several participants expressed worries about becoming overly dependent on ChatGPT (P1, P3, P4, P8, P11-P13). P1 cautioned about the “over-reliance on AI,” highlighting a tendency to “let it do it for you”. P6 and P7 also mirrored these concerns, indicating a habitual turn to ChatGPT as the first response to coding problems. P14 felt that ChatGPT was “most helpful if I already understood the concept, but just didn’t really want to spend the time implementing it myself.”, implying a trade-off between efficiency and educational depth.
@@
V-D5 - Marginal Role in Engagement & Curiosity Enhancement